Here's my problem:
Dumpster #1 contains 45 fewer molded vegetables than dumpster #2. Dumpster #2 contains 80 fewer molded vegetables than dumpster #3. If there are 236 molded vegetables in all the dumpsters combined, how many vegetables are in dumpster #1?

Answer by jorel555(1290)

Let n=dumpster 3. Then dumpster 2=n-80, and dumpster 1=n-80-45=n-125. So:
n+n-80+n-125=236
3n=441
n=147
n-80=67
n-125=22 ☺☺☺☺